### v3

# TaylorPODA: A Taylor Expansion-Based Method to Improve Post-Hoc Attributions for Opaque Models

Abstract
Existing post-hoc model-agnostic methods generate external explanations for opaque models, primarily by locally attributing the model output to its input features. However, they often lack an explicit and systematic framework for quantifying the contribution of individual features. Building on the Taylor expansion framework introduced by Deng et al. (2024) to unify existing local attribution methods, we propose a rigorous set of postulates — “precision”, “federation”, and “zero-discrepancy” — to govern Taylor term-specific attribution. Guided by these postulates, we introduce TaylorPODA (Taylor expansion-derived imPortance-Order aDapted Attribution), which incorporates an additional “adaptation” property. This property enables alignment with task-specific goals, especially in post-hoc settings lacking ground-truth explanations. Empirical evaluations demonstrate that TaylorPODA achieves competitive results against baseline methods, providing principled and visualization-friendly explanations. This work enhances the trustworthy deployment of opaque models by offering explanations with stronger theoretical grounding.
Introduction
The explainability of AI is becoming increasingly critical, particularly when users interact with models solely through an input-output interface, with only limited validation evidence inferred indirectly from test samples. This difficulty persists even when the model’s internal architecture and specific parameters are accessible, as the inherent opacity of many widely adopted models (e.g., deep neural networks) renders their underlying prediction mechanisms difficult for humans to interpret. Given the prevalence of such opaque models—whether due to restricted accessibility or intrinsic incomprehensibility—there is a growing need for external, model-agnostic methods to enhance their post-hoc explainability.
Local attribution (LA), as one of the predominant post-hoc strategies in explainable AI (XAI), focuses on allocating the model output value to each input feature. Unlike methods that provide global or average importance scores (Fisher, Rudin, and Dominici 2019; Gregorutti, Michel, and Saint-Pierre 2017), LA emphasizes the understanding of the contribution of each feature within a specific input in producing the particular output. Various LA methods, including LIME (Ribeiro, Singh, and Guestrin 2016) and SHAP (Lundberg and Lee 2017), have been developed and widely used in research and in practical applications. Although some LA methods, such as Integrated Gradients (Sundararajan, Taly, and Yan 2017) and TreeSHAP (Lundberg et al. 2020), require access to the model’s internal structure and parameters, most are model-agnostic, relying solely on its proposed approach to systematically query the model and compute feature-wise effects based on the outputs.
Through various designs, these post-hoc model-agnostic LA methods aim to marginalize and quantify the independent contribution of each feature. To unify the attribution paradigms underlying various methods, Deng et al. (2024) proposed an analytical framework based on Taylor expansion, assuming that the model output is a differentiable function of the input. From this perspective, the accuracy of feature attributions in existing post-hoc model-agnostic LA methods is undermined by two principal issues: (F1) the erroneous assignment of irrelevant Taylor terms to the target feature, and (F2) the imprecise distribution of Taylor terms, characterized by both incomplete utilization and overlapping allocations.
In current LA methods, there is a lack of coherent schemes within the Taylor framework to properly address (F1) and (F2). Additionally, interaction effects on the output are often oversimplified through fixed, pre-defined allocation — for example, SHAP assumes equal contributions among the involved features (see the Section “Analyzing existing methods through the lens of the Taylor framework”). Such allocations can potentially lead to arbitrary outcomes, deviating from the feature importance order of the instance under analysis. Moreover, given the limited knowledge of opaque models, especially in post-hoc contexts, such fixed and pre-defined assumptions further undermine the trustworthiness of the attribution outcomes. Therefore, there is a need for a more principled LA method with a flexible attribution process — one that is supported by rigorous analysis within the Taylor framework, as illustrated in Figure 1.
Contributions
This paper makes the following key contributions:
Figure 1: Illustration of the Taylor expansion framework used to analyze LA methods. While independent effects shall be allocated in a targeted manner, interaction effects require more careful handling, particularly in model-agnostic settings with limited prior knowledge. Many existing methods rely on fixed, pre-defined allocations, e.g., SHAP adopts a uniform distribution among the features involved.
  • A principled post-hoc model-agnostic LA framework. We introduce three postulates — precision, federation, and zero-discrepancy — under the Taylor expansion framework. These postulates formalize desiderata for accurately allocating Taylor terms in post-hoc model-agnostic attribution, effectively resolving (F1) the inclusion of non-relevant terms and (F2) the underuse or overlap of relevant terms.
  • A new explanation-providing method, TaylorPODA. We propose TaylorPODA that satisfies all the proposed postulates under the Taylor framework. More than that, TaylorPODA introduces an additional adaptation property by flexibly allocating Taylor interaction effects, enabling task-specific alignment via optimization objectives.
  • An AUP-driven optimization strategy. We instantiate the adaptation property by using the area under the prediction recovery curve (AUP) as a task-specific objective. With Dirichlet-distribution-based random search, TaylorPODA adaptively reallocates interaction terms to minimize AUP. This yields faithful attributions without relying on predefined, arbitrary prior assumptions.
  • Comprehensive empirical validation. We conduct extensive experiments on both tabular and image datasets, encompassing classification and regression tasks, and evaluate our method across various models. The evaluation comprises both quantitative analyses and qualitative visualizations, enabling a comprehensive assessment of attribution performance. Results show that TaylorPODA consistently matches or outperforms existing baselines, offering a more principled and trustworthy explanation-providing option under the Taylor framework. (Code is provided in the supplementary material)
Problem Formulation
In this section, we outline the foundational setup for this paper, starting with a description of how the Taylor expansion enables us to establish a systematic framework for understanding a LA method. Under this framework, several prominent methods are examined, including the widely-adopted Occlusion Sensitivity (Zeiler and Fergus 2014), SHAP (Lundberg and Lee 2017), and WeightedSHAP (Kwon and Zou 2022). Moreover, we include LIME (Ribeiro, Singh, and Guestrin 2016), a widely recognized post-hoc, model-agnostic method that, while not decomposable under the Taylor framework, also produces feature-wise importance scores. All of these methods are subsequently used as baselines in our experiments. We then introduce a set of postulates to establish a principled and consistent foundation for developing a more refined LA method that properly addresses (F1) and (F2).
Local attribution within the Taylor framework
Let $\mathcal{X}\subseteq\mathbb{R}^{d}$ be the input space for $d\in\mathbb{N}$, and let $\mathcal{Y}\subseteq\mathbb{R}$ be the output space. An opaque prediction model $f:\mathcal{X}\rightarrow\mathcal{Y}$ assigns an output $f(\mathbf{x})$ to an input $\mathbf{x}\in\mathcal{X}$ to produce the input-output instance pair $(\mathbf{x},f(\mathbf{x}))$. We assume that $f(\mathbf{x})$ is differentiable, following the settings in (Deng et al. 2024). Let $x_{1},...,x_{d}$ denote the exact feature values of $\mathbf{x}$ (indexed by $G=\{1,\dots,d\}$), then the $K$-order Taylor expansion of $f(\mathbf{x})$ at a baseline point $\boldsymbol{\beta}=(\beta_{1},...,\beta_{d})$ can be obtained as follows:
$f(\mathbf{x})=f(\boldsymbol{\beta})+\sum_{i\in G}\frac{1}{1!}\cdot\frac{\partial f(\boldsymbol{\beta})}{\partial x_{i}}\cdot(x_{i}-\beta_{i})$ (1)
$+\sum_{i\in G}\sum_{j\in G}\frac{1}{2!}\cdot\frac{\partial^{2}f(\boldsymbol{\beta})}{\partial x_{i}\partial x_{j}}\cdot(x_{i}-\beta_{i})(x_{j}-\beta_{j})+\dots+\epsilon_{K},$
where $\epsilon_{K}$ denotes the approximation error of Taylor expansion. This expansion unveils a dependence structure between features that opens the door to the grouping of the additive terms in (1) into Taylor independent effect and Taylor interaction effect.
Definition 1 (Taylor independent effect).
The additive terms in (1) which only involve one feature $i\in G$ are defined as Taylor independent effect:
$\lambda$ $(\boldsymbol{\phi})=\frac{1}{\phi_{i}!}\frac{\partial^{\phi_{i}}f(\boldsymbol{\beta})}{\partial x_{i}^{\phi_{i}}}(x_{i}-\beta_{i})^{\phi_{i}},$ (2)
where $\phi_{i}\in\mathbb{N}$ for $i\in G$, corresponding to the derivative order upon $x_{i}$, and $\boldsymbol{\phi}=(\phi_{1},\dots,\phi_{d})$ for $j\in G$ with $\phi_{j}=0$ when $j\neq i$.
Definition 2 (Taylor interaction effect).
The additive terms in (1) which involve more than one feature are identified as Taylor interaction effect:
$\mu$ $(\boldsymbol{\psi})=\frac{1}{(\psi_{1}+...+\psi_{d})!}\binom{\psi_{1}+...+\psi_{d}}{\psi_{1},...,\psi_{d}}$ (3)
$\cdot\frac{\partial^{\psi_{1}+...+\psi_{d}f}f(\boldsymbol{\beta})}{\partial x_{1}^{\psi_{1}}\dots\partial x_{d}^{\psi_{d}}}\cdot(x_{1}-\beta_{1})^{\psi_{1}}\dots(x_{d}-\beta_{d})^{\psi_{d}},$
where $\psi_{j}\in\mathbb{N}$ for $j\in G$ denoting the derivative order for $x_{j}$, and $\boldsymbol{\psi}=(\psi_{1},...,\psi_{d})$ with $|\{j\in G:\psi_{j}\neq 0\}|\geq 2$.
With Definition 1 and 2, $f(\mathbf{x})$ can be re-organized as follows:
$f(\mathbf{x})=f(\boldsymbol{\beta})+\sum_{i\in G}\sum_{\boldsymbol{\phi}\in\Pi_{\{i\}}}\lambda(\boldsymbol{\phi})+\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\end{subarray}}\sum_{\boldsymbol{\psi}\in\Pi_{S}}\mu(\boldsymbol{\psi}),$ (4)
where $\Pi_{T}=\{(\pi_{1},...,\pi_{d})\in\mathbb{N}^{d}|\pi_{j}=0\text{ for all }j\notin T\}$. Thus, $\Pi_{\{i\}}$ corresponds to all terms that are only partially derivative with respect to $x_{i}$, and $\Pi_{S}$ goes the similar way.
Based on Definition 1 and 2, LA methods can be generalized with the following form under this Taylor expansion framework.
Definition 3 (Local attribution).
Given a to-be-explained input-output pair $(\mathbf{x},f(\mathbf{x}))$, local attribution generates a group of contribution scores $\boldsymbol{a}=(a_{1},\dots,a_{d})$ with $a_{i}\in\mathbb{R}$ for $i\in G$, where the component $a_{i}$ measures the contribution of the corresponding $x_{i}$ by linearly combining the Taylor independent effects and the Taylor interaction effects within $f(\mathbf{x})$:
$a_{i}(\mathbf{x},f(\mathbf{x}))=\sum_{j\in G}\sum_{\boldsymbol{\phi}\in\Pi_{\{j\}}}\tau_{i,j}\lambda(\boldsymbol{\phi})+\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\end{subarray}}\sum_{\boldsymbol{\psi}\in\Pi_{S}}\zeta_{i,\psi}\mu(\boldsymbol{\psi}),$ (5)
where $\tau_{i,j},\zeta_{i,\psi}\in\mathbb{R}$. The weight $\tau_{i,j}$ quantifies the proportion of the Taylor independent effect $\lambda(\boldsymbol{\phi})$ from the $j$-th feature attributed to $x_{i}$. Similarly, the weight $\zeta_{i,\psi}$ represents the proportion of the Taylor interaction effect $\mu(\boldsymbol{\psi})$ from the features in $S$ attributed to $x_{i}$.
Before moving on to the next topic, we need to specify the masked output of the model $f$, since it has been used in the calculation process of most existing LA methods:
Definition 4 (Masked output).
Given an input $\mathbf{x}$ and a prediction model $f$, the corresponding masked output $f(\mathbf{x}_{S})$ is an estimated output to approximate a theoretic output of $f$ with the presence of the features in $S$, while eliminating the effect brought by the features outside $S$:
$f_{S}(\mathbf{x})=\mathbb{E}\left[f(x_{S},X_{G\backslash S})\right],$ (6)
where $(x_{S},X_{G\backslash S})=(X_{S_{1}},\dots,X_{S_{d}})$, with a slight abuse of notation.
Analyzing existing methods through the lens of the Taylor framework
Deng et al. (2024) explored several well-regarded LA methods by leveraging the Taylor framework outlined in (5). For example, Shapley value-based attribution has been demonstrated to be decomposable as follows (Given that the widely-adopted SHAP (Lundberg and Lee 2017) is fundamentally based on the Shapley value, we use “SHAP” to refer to any Shapley value-based post-hoc method in this paper, in order to avoid intricacy.):
$\underset{(\text{SHAP})}{a_{i}(\mathbf{x},f(\mathbf{x}))}$ $=\sum_{S\subseteq G\backslash\{i\}}p(S)\cdot\left[f_{S\cup\{i\}}(\mathbf{x})-f_{S}(\mathbf{x})\right]$ (7)
$=\sum_{\boldsymbol{\phi}\in\Pi_{\{i\}}}\lambda(\boldsymbol{\phi})+\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}\sum_{\boldsymbol{\psi}\in\Pi_{S}}\frac{1}{|S|}\mu(\boldsymbol{\psi}),$
where p(S)=|S|!(|G|1|S|)!/|G|!p(S)=|S|!(|G|-1-|S|)!/|G|!. And for another attribution method, Occlusion Sensitivity (referred to as OCC-1 below, named in accordance with (Deng et al. 2024)), has been shown the following decomposition:
[…]
$=\sum_{\boldsymbol{\phi}\in\Pi_{\{i\}}}\lambda(\boldsymbol{\phi})+\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}\sum_{\boldsymbol{\psi}\in\Pi_{S}}\mu(\boldsymbol{\psi}).$
All of the methods analyzed in (Deng et al. 2024) adopted a fixed allocation of the Taylor terms, whereas WeightedSHAP (Kwon and Zou 2022), one of the inspiring variants of SHAP, started to flexibly and adaptively provide post-hoc explanations. Here, we use the same Taylor expansion-based LA framework to investigate WeightedSHAP:
[…]
$=\sum_{S\subseteq G\backslash\{i\}}w_{S}\left[\sum_{\boldsymbol{\phi}\in\Pi_{\{i\}}}\lambda(\boldsymbol{\phi})+\sum_{\begin{subarray}{c}i\in T\\ T\backslash\{i\}\subseteq S\end{subarray}}\sum_{\boldsymbol{\psi}\in\Pi_{T}}\mu(\boldsymbol{\psi})\right],$
where $w_{S}\in\mathbb{R^{+}}$, used as an adaptive weight.
Although not an attribution-based method, we include the widely-adopted LIME (Ribeiro, Singh, and Guestrin 2016) to facilitate clearer analysis:
$\underset{(\text{LIME})}{g(\mathbf{x},f(\mathbf{x}))}=\sum_{i\in G}\eta_{i}\cdot x_{i}\approx f(\mathbf{x}),$ (10)
where $\eta_{i}\in\mathbb{R}$ for $i\in G$, representing the feature contribution, and $\boldsymbol{\eta}=(\eta_{1},\dots,\eta_{d})$ is obtained as follows:
$\boldsymbol{\eta}(\mathbf{x},f(\mathbf{x}))=\underset{g\in\Lambda}{\arg\min}\mathcal{L}(f,g,\iota_{\mathbf{x}})+\theta(g),$ (11)
where $\Lambda$ represents the explanation family $\boldsymbol{\eta}\subseteq\mathbb{R}^{d}$, $\iota_{\mathbf{x}}$ defines the weight for the distance function $\mathcal{L}$, and $\theta(g)$ regulates the complexity of $g$.
Most of the existing post-hoc methods that provide model-agnostic explanations with feature-wise importance scores can be defined as a variant or an approximation of the methods mentioned above, especially the large family of Shapley value-based ones (Štrumbelj and Kononenko 2014; Aas, Jullum, and Løland 2021; Albini et al. 2022; Jethani et al. 2022; Kolpaczki et al. 2024).
Absent postulate Example Taylor term Potential consequences
Precision $\frac{1}{5!}\cdot\frac{\partial^{5}f(\boldsymbol{\beta})}{\partial x_{2}^{5}}\cdot(x_{2}-\beta_{2})^{5}$
The Taylor term can be attributed to $x_{1}$ and
included in $a_{1}$, rather than being fully attributed
to $x_{2}$ and included in $a_{2}$.
Federation $\frac{1}{2!}\cdot\frac{\partial^{2}f(\boldsymbol{\beta})}{\partial x_{1}\partial x_{2}}\cdot(x_{1}-\beta_{1})\cdot(x_{2}-\beta_{2})$
The Taylor term can be attributed to $x_{3}$ and
included in $a_{3}$.
Zero
-discrepancy
$\frac{1}{3!}\cdot\frac{\partial^{3}f(\boldsymbol{\beta})}{\partial x_{1}\partial x_{2}\partial x_{3}}\cdot(x_{1}-\beta_{1})\cdot(x_{2}-\beta_{2})\cdot(x_{3}-\beta_{3})$
An incomplete allocation of the Taylor term
(e.g., only 90% distributed among $a_{1},a_{2},a_{3}$),
which leads to $f(\boldsymbol{\beta})+a_{1}+a_{2}+a_{3}\neq f(\mathbf{x})$.
Table 1: Illustrative examples of postulate omission consequences.
Postulates towards a principled LA method under the Taylor framework
Under the Taylor framework discussed above, we introduce the following postulates to further refine and regulate the LA defined with (3), in order to better address (F1) and (F2):
Postulate 1.
Precision. The Taylor independent effect of the $i$-th feature shall be entirely attributed to the $i$-th feature, while it shall not be attributed to any other feature:
[…]
Postulate 2.
Federation. The Taylor interaction effect of the features in $S$ shall only be attributed to the features inside $S$:
$\zeta_{i,\psi}=0,\quad\text{for all }i\notin S,\ \boldsymbol{\psi}\in\Pi_{S}.$ (13)
Postulate 3.
Zero-discrepancy. There should be neither redundancy nor deficiency in the attribution results regarding the allocation of the exact model output $f(\mathbf{x})$ to individual features. Equivalently, the value of discrepancy, denoted by $d(\mathbf{x},f;\boldsymbol{\beta};\mathbf{a})$ shall equal zero:
$d(\mathbf{x},f;\boldsymbol{\beta};\mathbf{a})=f(\boldsymbol{\beta})+\sum_{i\in G}a_{i}(\mathbf{x},f(\mathbf{x}))-f(\mathbf{x})=0.$ (14)
This postulate aligns with the design principle of “local accuracy(Lundberg and Lee 2017), a core property of SHAP derived from the “efficiency” axiom of the Shapley value, where “the value represents a distribution of the full yield of the game” (Shapley 1953).
Among the proposed postulates, precision and federation address (F1) by including all Taylor terms involving the $i$-th feature while excluding unrelated terms. Zero-discrepancy addresses (F2) by ensuring that the sum of attributions exactly matches the model output, fully allocating it across relevant features.
To illustrate the necessity of these postulates, Table 1 presents examples within the Taylor framework (under a three-feature setting, $G=\{1,2,3\}$). These cases show how omitting any single postulate may lead to inconsistent or counterintuitive attribution results.
Proposed Methodology
We propose a new LA method, TaylorPODA, which fulfills all postulates 1, 2, and 3. In addition, TaylorPODA introduces a desirable property—adaptation (see property 1)—by formulating the attribution of Taylor terms associated with interaction effects as an optimizable process guided by a user-defined objective (with AUP adopted in this work), thereby aligning the inherent ambiguity of ground-truth explanations in post-hoc contexts with the downstream task-specific target.
Taylor expansion-derived importance-order adapted attribution
Given an input-output pair $(\mathbf{x},f(\mathbf{x}))$, TaylorPODA attributes the output $f(\mathbf{x})$ to the $i$-th feature as follows:
[…]
where the Harsanyi dividend $H(S)=\sum_{T\subseteq S}(-1)^{|T|-|S|}f_{T}(\mathbf{x})$, originally proposed in a game-theoretic context by Harsanyi (1982), is here applied as an operator over masked model outputs, and the coefficients $\xi_{i,S}\in(0,1)$ are tunable weights introduced to enable further adaptation based on the importance ordering, subject to the constraint $\sum_{i\in S}\xi_{i,S}=1$ for any subset $S\subseteq G$ with $|S|>1$.
Postulate and property satisfaction: TaylorPODA and other methods
The attribution value produced by TaylorPODA, as defined in (15), is equivalent to the following Taylor-term representation, which consists of Taylor independent effects and Taylor interaction effects:
$\underset{(\text{TaylorPODA})}{a_{i}(\mathbf{x},f(\mathbf{x}))}=\sum_{j=i}\sum_{\boldsymbol{\phi}\in\Pi_{\{j\}}}\lambda(\boldsymbol{\phi})+\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}\sum_{\boldsymbol{\psi}\in\Pi_{S}}\xi_{i,S}\mu(\boldsymbol{\psi}).$ (16)
Moreover, inspired by (Kwon and Zou 2022)—which highlights that “the Shapley value incorrectly reflects the influence of features, resulting in a suboptimal order of attributions”—TaylorPODA introduces tunable coefficients $\xi_{i,S}\in(0,1)$ as allocation weights $\zeta_{i,S}$ in the LA structure (5), in contrast to SHAP’s pre-defined uniform allocation. This flexibility enables optimizable attribution and gives rise to an advantageous property—adaptation—as defined below:
Property 1.
Adaptation. For the $i$-th feature, the proportion of attribution from each Taylor interaction effect $\mu(\boldsymbol{\psi})$ with $\boldsymbol{\psi}\in\Pi_{S}$ and $S\subseteq G\setminus\{i\},|S|>1$ shall be tunable. Specifically, the attribution mechanism allows $\zeta_{i,\boldsymbol{\psi}}\in\left[0,1\right]$ for all $\boldsymbol{\psi}\in\Pi_{S}$ with $S\subseteq G,|S|>1,i\in S$.
This property enables the attribution mechanism to flexibly allocate Taylor interaction effects among involved features based on task-specific optimization objectives, particularly in post hoc, model-agnostic settings where ground-truth explanations are typically unavailable with an opaque model.
To the best of our knowledge, no existing post-hoc model-agnostic method—including OCC-1, LIME, SHAP, and WeightedSHAP—satisfies all these postulates and properties, as shown in Table 2.
Methods PRC FDR ZDC ADT
OCC-1 $\checkmark$ $\checkmark$ $\times$ $\times$
LIME
SHAP $\checkmark$ $\checkmark$ $\checkmark$ $\times$
WeightedSHAP $\times$ $\checkmark$ $\times$ $\checkmark$
TaylorPODA $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 2: Postulate and property satisfaction. TaylorPODA is the only method that satisfies all the postulates and property. (PRC for precision, FDR for federation, ZDC for zero-discrepancy, ADT for adaptation.)
Classification Regression
Method
Data AUP Inclusion AUC Data AUP Inclusion MSE ($\times 10^{-2}$)
OCC-1
Cancer
0.672 (0.624, 0.720) 0.996 (0.991, 1.000)
Abalone
0.152 (0.133, 0.171) 0.076 (0.057, 0.096)
LIME 0.790 (0.736, 0.844) 0.991 (0.982, 1.000) 0.140 (0.124, 0.156) 0.052 (0.038, 0.066)
SHAP 0.874 (0.792, 0.956) 0.981 (0.962, 1.000) 0.161 (0.143, 0.178) 0.062 (0.048, 0.076)
WeightedSHAP 0.519 (0.483, 0.555) 0.998 (0.995, 1.000) 0.104 (0.090, 0.117) 0.037 (0.028, 0.047)
TaylorPODA 0.601 (0.530, 0.673) 0.991 (0.983, 1.000) 0.092 (0.081, 0.103) 0.026 (0.020, 0.031)
OCC-1
Rice
0.595 (0.526, 0.665) 0.986 (0.968, 1.000)
California
0.171 (0.147, 0.195) 0.133 (0.099, 0.167)
LIME 0.694 (0.611, 0.776) 0.986 (0.968, 1.000) 0.263 (0.232, 0.294) 0.296 (0.221, 0.371)
SHAP 0.668 (0.574, 0.763) 0.990 (0.976, 1.000) 0.186 (0.162, 0.210) 0.135 (0.101, 0.170)
WeightedSHAP 0.470 (0.408, 0.531) 0.991 (0.980, 1.000) 0.133 (0.114, 0.153) 0.091 (0.065, 0.117)
TaylorPODA 0.493 (0.427, 0.559) 0.991 (0.975, 1.000) 0.154 (0.132, 0.176) 0.108 (0.077, 0.139)
OCC-1
Titanic
0.530 (0.470, 0.591) 0.961 (0.937, 0.986)
Concrete
0.373 (0.334, 0.411) 0.486 (0.384, 0.587)
LIME 0.625 (0.552, 0.699) 0.946 (0.913, 0.979) 0.343 (0.313, 0.372) 0.365 (0.305, 0.426)
SHAP 0.516 (0.461, 0.571) 0.969 (0.943, 0.995) 0.274 (0.251, 0.296) 0.253 (0.213, 0.293)
WeightedSHAP 0.392 (0.345, 0.439) 0.964 (0.941, 0.987) 0.226 (0.204, 0.248) 0.197 (0.158, 0.235)
TaylorPODA 0.444 (0.392, 0.496) 0.973 (0.952, 0.994) 0.221 (0.199, 0.244) 0.193 (0.154, 0.231)
Table 3: Post-hoc method performance on 100 hold-out test samples per dataset in reflecting the feature importance ordering measured by AUP, Inclusion AUC, and Inclusion MSE with means ± 95% confidence intervals. (AUP and Inclusion MSE: lower–better. Inclusion AUC: higher–better.)
Proof: The proof of (15)’s equivalence to (16), along with its compliance with Postulates 1, 2, and 3, is provided in Appendix A, in which a detailed analysis of how these postulates and property are violated by the other methods is also included. $\blacksquare$
Optimization strategy
To obtain an optimal $\xi_{i,S}$ in (15) when using TaylorPODA, we adopt a metric, area under the prediction recovery error curve (AUP) introduced by WeightedSHAP (Kwon and Zou 2022), as the optimization objective in TaylorPODA. AUP serves to evaluate the feature-importance-order alignment degree between the absolute attribution values $|a_{i}|$ and the model prediction on $\mathbf{x}$. Specifically, given a group of attribution scores $\mathbf{a}=(a_{1},...,a_{d})$ for each feature within $(\mathbf{x},f(\mathbf{x}))$, let $\mathcal{I}(m;\mathbf{a},\mathbf{x})\subseteq\left[d\right]$ be a set of $m$ integers that indicates the top-$m$ most important features in terms of $|a_{i}|$. With $\mathcal{I}(m;\mathbf{a},\mathbf{x})$, we have AUP as follows:
$\text{AUP}(\mathbf{a};\mathbf{x},f)$ (17)
$=\sum_{m=1}^{d}\bigg{|}f(\mathbf{x})-\mathbb{E}\left[f(X)|X_{\mathcal{I}(m;\mathbf{a},\mathbf{x})}=\mathbf{x}_{\mathcal{I}(m;\mathbf{a},\mathbf{x})}\right]\bigg{|}.$
In this version of TaylorPODA, the coefficients $\xi_{i,S}$ are determined by solving an optimization problem with AUP as the objective:
$\boldsymbol{\xi}^{*}$ $=\underset{\boldsymbol{\xi}\in\boldsymbol{Z}}{\arg\min}\ \text{AUP}(\boldsymbol{a};\mathbf{x},f)$ (18)
$\text{s.t.}\quad\sum_{i\in S}\xi_{i}$ $=1,\quad\forall S\subseteq G\text{ with }|S|>1,$
where $\boldsymbol{Z}$ represents the coefficient family encompassing all possible coefficient vectors $\boldsymbol{\xi}=(\xi_{S_{1},1},\dots,\xi_{S_{1},d},\dots,\xi_{S_{q},1},\dots,\xi_{S_{q},d})$ with $q=2^{d-1}$ for all $S\subseteq G$, $|S|>1$. The AUP-based optimization in (18) serves to illustrate how the attribution process can be adapted to instance-specific faithfulness criteria. The choice of optimization objective is task-dependent and user-configurable based on downstream needs.
Furthermore, this work adopts a random search approach to obtain a solution of (18) for TaylorPODA. Based on Dirichlet distribution, a set of candidate solutions that satisfy the constraints in (18) are produced. From these candidates, the solution achieving the least AUP is selected as $\boldsymbol{\xi}^{*}$. Formally, given a feature coalition $S$ and a concentration parameter vector $\boldsymbol{\alpha_{S}}=(\alpha_{r_{1},S},\dots,\alpha_{r_{|S|},S})$, the Dirichlet distribution samples a vector $\boldsymbol{\xi}_{S}=(\xi_{r_{1},S},\dots,\xi_{r_{|S|},S})$ for the features $r_{1},\dots,r_{|S|}\in S$, where $\xi_{r_{i},S}\in(0,1)$ for $i\in S$. The probability density function of Dirichlet distribution is given by:
$p(\boldsymbol{\xi}_{S};\boldsymbol{\alpha}_{S})=\frac{\Gamma\left(\sum_{i=1}^{|S|}\alpha_{i,S}\right)}{\prod_{i=1}^{|S|}\Gamma(\alpha_{i,S})}\prod_{i=1}^{|S|}\xi_{r_{i},S}^{\alpha_{i,S}-1},$ (19)
where $\Gamma(\cdot)$ is the Gamma function, and the vector $\boldsymbol{\alpha}_{S}$ is set to control the extent to which the generated $\boldsymbol{\xi}_{S}$ deviates from the uniform distribution (equivalent to a Shapley value result). Regulated by the properties of the Dirichlet distribution, the generated $\boldsymbol{\xi}_{S}$ based on (19) inherently satisfies the requirement $\sum_{i=1}^{|S|}\xi_{r_{i},S}=1$ in (18). This ensures that Postulate 3 (zero-discrepancy) is met, as demonstrated in Appendix A. See (Ng, Tian, and Tang 2011) for the further details and properties of Dirichlet distribution.
Experimental Results
The effectiveness of the proposed TaylorPODA is evaluated on several datasets covering both regression and classification tasks. Widely-recognized post-hoc model-agnostic attribution schemes represented by OCC-1, LIME, SHAP, and WeightedSHAP, are used as the baselines. Details of the datasets and the implementation specifics are provided in Appendix B.
Feature importance alignment.
We evaluate the efficiency of TaylorPODA and baseline methods in capturing correct feature importance orderings. Since both TaylorPODA and WeightedSHAP optimize AUP directly, we also report two complementary metrics to enhance comparability: Inclusion AUC (Jethani et al. 2022) for classification and its regression counterpart, Inclusion MSE (Kwon and Zou 2022). All three metrics assess the quality of importance orderings derived from absolute attribution values.
Table 3 summarizes results on differentiable opaque models with multilayer perceptron structures using tanh and logistic activations. Across all metrics, TaylorPODA and WeightedSHAP consistently alternate in top performance among various datasets, indicating their strong alignment with instance-level feature importance—not only in minimizing AUP but also in achieving high Inclusion AUC and low Inclusion MSE.
Although the experimental settings in Table 3 comply with the differentiability requirement of Taylor expansion, the TaylorPODA formulation in (15) remains applicable to non-differentiable models—similar to Shapley value-based methods, which are practically used without regard to model differentiability, even though the Taylor framework assumes it for theoretical interpretation. Empirically, experiments on non-differentiable models also yield results consistent with the conclusions of this section (see Appendix C).
Figure 2: Violin plots of discrepancy performance across 100 hold-out test samples per dataset. Both SHAP and TaylorPODA consistently exhibited zero discrepancy.
Figure 3: SHAP vs. TaylorPODA explanations for the same sample from the Concrete dataset (Yeh 1998). Both satisfy the zero-discrepancy property, enabling such bar-plot visualization. TaylorPODA yields a lower (better) AUP.
Discrepancy alignment.
The absolute discrepancy measures the value gap between attribution outcomes and the corresponding model prediction. As shown in the violin plots (Figure 2), both TaylorPODA and SHAP consistently satisfy the zero-discrepancy postulate across all test samples, aligning with the theoretical analysis in the previous section. These results reflect that both methods exhaustively allocate Taylor expansion terms without redundancy or omission. In contrast, other methods violate the zero-discrepancy postulate, showing non-zero discrepancies—positive values indicating overlapping attribution, and negative values implying under-allocation. Moreover, as illustrated in Figure 3, TaylorPODA supports SHAP-style visualization by providing feature-wise contributions for individual predictions. This is because of its satisfaction of the zero-discrepancy postulate, which allows the summarization of the feature-wise scores to align with the particular model output.
Qualitative and visualization alignment.
Qualitatively, TaylorPODA  [^1] Here, we adopt a heuristic approximation of the fully enumerative version of TaylorPODA in (15) by imposing an upper bound $c$ on the cardinality of feature subsets, restricting to $|S|\leq c$ instead of exhaustively traversing all $S\subseteq G$. Otherwise, with the $28\times 28$ MNIST image input, the full version of TaylorPODA defined in (15) would require computing $H(S)$ for $2^{28\times 28-1}$ subsets, which is computationally infeasible. Further analyses of this approximation and its error are provided in Appendix D. demonstrates competitive performance on image data. As illustrated in Figure 4 based on MNIST (LeCun et al. 1998), the attribution patterns generated by TaylorPODA exhibit high consistency with those produced by SHAP and WeightedSHAP, while also aligning well with visual intuition. In particular, TaylorPODA consistently highlights the openness of the left segments of the upper and lower loops of digit “8” as key discriminative features—regions that effectively distinguish it from digit “3”.
Figure 4: Illustrative examples of attributing model predictions for classifying digit-3 and digit-8 in MNIST. The pixels are color-coded: blue for negative, red for positive contribution to predicting “8”.
Related Work
Significant efforts have been made to enhance the explainability of opaque models. Early methods, such as partial dependence plots (Friedman 2001) and individual conditional expectation (Goldstein et al. 2015), deeply incorporate visualizing the fluctuations in the model output by altering the feature values. Later, more specific post-hoc methods emerged. Ribeiro, Singh, and Guestrin (2016) proposed LIME, which builds an interpretable local surrogate for the original model. In parallel, LA methods were developed to marginalize the contribution of individual features. OCC-1 (Zeiler and Fergus 2014), based on prediction difference (Robnik-Šikonja and Kononenko 2008), computes LA by masking target features. Štrumbelj and Kononenko (2014) extended Shapley value (Shapley 1953) to opaque models in AI context. Lundberg and Lee (2017) introduced SHAP and its popular implementation, which inspired many variants (Frye, Rowat, and Feige 2020; Aas, Jullum, and Løland 2021; Watson 2022). Among them, WeightedSHAP (Kwon and Zou 2022) relaxes SHAP’s “local accuracy” to adopt a more flexible semi-value (Dubey and Weber 1977; Hart and Mas-Colell 1989), and uses AUP to better align attributions with instance-level feature importance orderings.
Building on feature coalition-level attribution, Sundararajan, Dhamdhere, and Agarwal (2020) proposed the Shapley-Taylor Interaction Index, assigning scores to feature subsets and relating them to Taylor expansion terms. While this captures interactions, it introduces scalability and interpretability issues as subset numbers grow exponentially. To focus on individual feature-level attributions, Deng et al. (2024) proposed a Taylor expansion framework that unifies various post-hoc methods, including Shapley-based ones.
Conclusion and Future Work
We propose TaylorPODA, a new post-hoc model-agnostic method for LA that quantifies feature-wise contributions. It produces explanations that better align with feature importance while ensuring exact and exhaustive allocation of the model output. Also, TaylorPODA enables “SHAP-like” visualizations, offering readily understandable explainability. Furthermore, this method supports a user-configurable optimization process, allowing downstream, implementation-specific objectives to be flexibly incorporated. Significantly, the underlying postulates and property reinforce the theoretical foundation for trustworthy deployment of TaylorPODA and help avoid the crucial yet often overlooked paradox of explaining opacity with opacity.
Nonetheless, improving the computational efficiency of TaylorPODA remains an open challenge. As defined in (15), full evaluation requires computing $2^{|G|-1}$ Harsanyi dividends, each involving $2^{|S|}$ masked output queries. Encouragingly, our experiments demonstrate that truncating $|S|$ still yields results comparable to other post-hoc baselines (noting that, in the model-agnostic setting, no definitive ground-truth explanation exists). Still, more advanced approximation strategies are needed to enhance scalability.
Acknowledgments
This work was supported by the UK Research and Innovation (UKRI) Engineering and Physical Sciences Research Council (EPSRC) Doctoral Training Partnership (DTP) through the Healthy Lifespan Institute (HELSI) Flagship Scholarship at the University of Sheffield (Grant No. EP/W524360/1).
Appendix A A. Proof of Postulate and Property Satisfaction: TaylorPODA and Other Methods
Given a to-be-explained input-output pair $(\mathbf{x},f(\mathbf{x}))$, under a Taylor-expansion framework, a LA generates a group of contribution scores $\boldsymbol{a}=(a_{1},\dots,a_{d})$ with $a_{i}\in\mathbb{R}$ for $i\in G$, where the component $a_{i}$ measures the contribution of the corresponding $x_{i}$ by linearly combining the Taylor independent effects and the Taylor interaction effects within $f(\mathbf{x})$:
$a_{i}(\mathbf{x},f(\mathbf{x}))=\sum_{j\in G}\sum_{\boldsymbol{\phi}\in\Pi_{\{j\}}}\tau_{i,j}\lambda(\boldsymbol{\phi})+\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\end{subarray}}\sum_{\boldsymbol{\psi}\in\Pi_{S}}\zeta_{i,\psi}\mu(\boldsymbol{\psi}),$ (20)
where $\tau_{i,j},\zeta_{i,\psi}\in\mathbb{R}$. The weight $\tau_{i,j}$ quantifies the proportion of the Taylor independent effect $\lambda(\boldsymbol{\phi})$ from the $j$-th feature attributed to $x_{i}$. Similarly, the weight $\zeta_{i,\psi}$ represents the proportion of the Taylor interaction effect $\mu(\boldsymbol{\psi})$ from the features in $S$ attributed to $x_{i}$.
TaylorPODA, as a LA, produces the attribution outcome for feature $i$ as follows. Let $\xi_{i,S}\in(0,1)$ for $S\subseteq G$ and $i\in S$, satisfying $\sum_{i\in S}\xi_{i,S}=1$:
[…]
which is equivalent to
$\underset{(\text{TaylorPODA})}{a_{i}(\mathbf{x},f(\mathbf{x}))}=\sum_{j=i}\sum_{\boldsymbol{\phi}\in\Pi_{\{j\}}}\lambda(\boldsymbol{\phi})+\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}\sum_{\boldsymbol{\psi}\in\Pi_{S}}\xi_{i,S}\mu(\boldsymbol{\psi}).$ (22)
Moreover, TaylorPODA meets Postulates 1, 2, 3, and Property 1, whereas other methods (OCC-1, LIME, SHAP, WeightedSHAP) fail to satisfy all of these postulates and the property collectively:
Postulate 1. Precision. The Taylor independent effect of the $i$-th feature shall be entirely attributed to the $i$-th feature, while it shall not be attributed to any other feature:
[…]
Postulate 2. Federation. The Taylor interaction effect of the features in $S$ shall only be attributed to the features inside $S$:
$\zeta_{i,\psi}=0,\quad\text{for all }i\notin S,\ \boldsymbol{\psi}\in\Pi_{S}.$ (24)
Postulate 3. Zero-discrepancy. There should be neither redundancy nor deficiency in the attribution results regarding the allocation of the exact model output $f(\mathbf{x})$ to individual features. Equivalently, the value of discrepancy, denoted by $d(\mathbf{x},f;\boldsymbol{\beta};\mathbf{a})$ shall equal zero:
$d(\mathbf{x},f;\boldsymbol{\beta};\mathbf{a})=f(\boldsymbol{\beta})+\sum_{i\in G}a_{i}(\mathbf{x},f(\mathbf{x}))-f(\mathbf{x})=0.$ (25)
Property 1. Adaptation. For the $i$-th feature, the proportion of attribution from each Taylor interaction effect $\mu(\boldsymbol{\psi})$ with $\boldsymbol{\psi}\in\Pi_{S}$ and $S\subseteq G\setminus\{i\},|S|>1$ shall be tunable. Specifically, the attribution mechanism allows $\zeta_{i,\boldsymbol{\psi}}\in\left[0,1\right]$ for all $\boldsymbol{\psi}\in\Pi_{S}$ with $S\subseteq G,|S|>1,i\in S$.
Proof:
According to Theorem 2 in (Deng et al. 2024), we have:
[…]
Also, according to Theorem 1 in (Deng et al. 2024), we have
$H(S)=\sum_{\boldsymbol{\psi}\in\Pi_{S}}\mu(\boldsymbol{\psi}),\forall S\in G,|S|>1.$ (27)
Substituting (26) and (27) into (21), we get
[…]
$=\sum_{j=i}\sum_{\boldsymbol{\phi}\in\Pi_{\{j\}}}\lambda(\boldsymbol{\phi})+\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}\sum_{\boldsymbol{\psi}\in\Pi_{S}}\mu(\boldsymbol{\psi})-\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}(1-\xi_{i,S})H(S)$
$=\sum_{j=i}\sum_{\boldsymbol{\phi}\in\Pi_{\{j\}}}\lambda(\boldsymbol{\phi})+\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}H(S)-\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}(1-\xi_{i,S})H(S)$
$=\sum_{j=i}\sum_{\boldsymbol{\phi}\in\Pi_{\{j\}}}\lambda(\boldsymbol{\phi})+\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}\sum_{\boldsymbol{\psi}\in\Pi_{S}}\xi_{i,S}\mu(\boldsymbol{\psi}).$
Thus, TaylorPODA defined in (21) is equivalent to (22). Moreover, by setting $\tau_{i,j}=1$ and $\zeta_{i,\psi}=\xi_{i,S}$ for $\psi\in\Pi_{S}$, (20) can be equivalently written as (22).
Therefore, when calculating $\underset{(\text{TaylorPODA})}{a_{i}(\mathbf{x},f(\mathbf{x}))}$ only the Taylor independent effect of the $i$-th feature, i.e., $\sum_{j=i}\sum_{\boldsymbol{\phi}\in\Pi_{\{j\}}}\lambda(\boldsymbol{\phi})$, are involved without any other features $j\neq i$, as demonstrated by (21). Thus, TaylorPODA satisfies Postulate 1. Similarly, according to (8) and (7), OCC-1 and SHAP satisfy Postulate 1. As demonstrated in (9), WeightedSHAP attributes Taylor independent effects with a weighting factor $w_{S}$, thereby violating Postulate 1.
As indicated in (22), the Taylor interaction effects will be attributed to the $i$-th feature, if and only if $S\subseteq G\$with $i\in S$ and $|S|>1$. Thus, TaylorPODA satisfies Postulate 2. Similarly, according to (8), (7), and (9), it can be found that OCC-1, SHAP, and WeightedSHAP satisfy Postulate 1.
As for discrepancy, we have
$f(\boldsymbol{\beta})+\sum_{i\in G}\underset{(\text{TaylorPODA})}{a_{i}(\mathbf{x},f(\mathbf{x}))}$ (29)
$=$ $f(\boldsymbol{\beta})+\sum_{i\in G}\left[\sum_{j=i}\sum_{\boldsymbol{\phi}\in\Pi_{\{j\}}}\lambda(\boldsymbol{\phi})+\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}\xi_{i,S}\sum_{\boldsymbol{\psi}\in\Pi_{S}}\mu(\boldsymbol{\psi})\right]$
$=$ $f(\boldsymbol{\beta})+\sum_{i\in G}\sum_{\boldsymbol{\phi}\in\Pi_{\{i\}}}\lambda(\boldsymbol{\phi})+\sum_{i\in G}\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}\xi_{i,S}\sum_{\boldsymbol{\psi}\in\Pi_{S}}\mu(\boldsymbol{\psi}).$
Given that $\sum_{i\in S}\xi_{i,S}=1$ for $S\subseteq G$, (29) can be further transformed into:
$f(\boldsymbol{\beta})+\sum_{i\in G}\underset{(\text{TaylorPODA})}{a_{i}(\mathbf{x},f(\mathbf{x}))}=f(\boldsymbol{\beta})+\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\end{subarray}}\sum_{\boldsymbol{\phi}\in\Pi_{\{i\}}}\lambda(\boldsymbol{\phi})+\sum_{i\in G}\sum_{\boldsymbol{\psi}\in\Pi_{S}}\mu(\boldsymbol{\psi})=f(\mathbf{x}).$ (30)
Thus, TaylorPODA satisfies Postulate 3. Similarly, as given in (8), we can equivalently have $\zeta_{i,S}=1$ for $i\in G$ and $S\subseteq G$ with $|S|>1$, so that $\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}\xi_{i,S}>1$ for OCC-1. Similarly again, as given in (7), we can equivalently have $\xi_{i,S}=1/|S|$ for $i\in G$ and $S\subseteq G$ with $|S|>1$, so that $\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}\xi_{i,S}=1$ for SHAP. However, as the weighting factor $\omega_{S}$ is not limited in terms of its sum value, it is not ensured that $\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}\xi_{i,S}=1$ in WeightedSHAP. Thus, SHAP satisfies Postulate 3, whereas OCC-1 and WeightedSHAP violate Postulate 3.
Moreover, as $\xi_{i,S}$ is an adaptive weight, essentially, the exact quantity of the Taylor interaction that is to be allocated to the $i$-th feature is adjustable, instead of setting a fixed ratio. Thus, TaylorPODA introduces (satisfies) Property 1. Similarly, according to (8), (7), and (9), it can be found that WeightedSHAP satisfies Property 1, whereas OCC-1 and SHAP violate Property 1.
As for LIME, it should not be regarded as a strict LA method or even an attributional method for allocating the contribution of each feature in $f(\mathbf{x})$, since it explains models by introducing an external surrogate model $g(\mathbf{x})$ to approximate the original model, as shown by (10) and (11). Consequently, LIME falls outside the scope of the postulate (property) system in this work, and its corresponding columns in Table 2 are marked with “–”.
This completes the proof. $\blacksquare$
Appendix B B. Implementation Details of the Experiments
All experiments were conducted on a machine equipped with a 13th Gen Intel® Core™ i7-13700K CPU (3.40 GHz) and 32 GB RAM.
The datasets and the corresponding prediction tasks:
Seven tabular datasets together with a two-dimensional image dataset are used for the experiments, all of which are publicly available. Based on these datasets, the corresponding prediction tasks are designed. Details of these datasets and the are shown as follows:
Table 4: Details of the datasets and the corresponding prediction tasks used for the experiments.
Dataset Size Features Task Source
Cancer 683 Clump_thickness, Uniformity_of_cell_size, Uniformity_of_cell_shape, Marginal_adhesion, Single_epithelial_cell_size, Bare_nuclei, Bland_chromatin, Normal_nucleoli, Mitoses Classifying whether the (Class) of the cancer is benign or malignant. https://archive.ics.uci.edu/dataset/15/breast+cancer+wisconsin+original
Rice 3810 Area, Perimeter, Major_Axis_Length, Eccentricity, Convex_Area, Extent Classifying whether the rice specie (Class) is Osmancik or Cammeo. https://archive.ics.uci.edu/dataset/545/rice+cammeo+and+osmancik
Titanic 712 Age, Fair, Pclass, Sibsp, Parch, Alone, Adult_male Classifying whether the passenger is survival or not. https://www.kaggle.com/c/titanic/data
Abalone 4177 Sex, Length, Height, Whole_weight, Shucked_weight, Viscera_weight, Shell_weight Predicting the age (rings) of abalone from physical measurements. https://archive.ics.uci.edu/dataset/1/abalone
California 20640 MedInc, HouseAge, AveRooms, AveBedrms, Population, AveOccup, Latitude, Longitude Predicting the median house value (MedHouseVal) of California districts from demographic and geographic information. https://scikit-learn.org/1.5/modules/generated/sklearn.datasets.fetch˙california˙housing.html
Concrete 1030 Cement, Blast furnace slag, Fly ash, Water, Superplasticizer, Coarse aggregate, Fine aggregate, Age Predicting the compressive strength of concrete mixtures from their ingredients and age. https://archive.ics.uci.edu/dataset/165/concrete+compressive+strength
MNIST38 13966 Two-dimensional grayscale image pixels Classifying the digits 3 and 8 from the hand-written numbers. https://keras.io/api/datasets/mnist/
All datasets are shuffled, with 80% of the samples in each set randomly selected for training. Attribution experiments are conducted using 100 hold-out samples randomly drawn from the remaining 20%.
The $28\times 28$ MNIST38 dataset is obtained from the original MNIST dataset by extracting all the digit-3 and digit-8 images while excluding the images of other digits.
The task models:
For the prediction tasks with all the datasets, we adopt machine learning-based fully connected multi-layer perceptron (MLP).
Regarding differentiability, it is important to note that all activation functions used in the quantitative analysis presented in the main body (specifically, the experiments related to Table 3)—namely, tanh and logistic—are continuously differentiable. As a result, the MLP-based task models are fully differentiable, enabling proper evaluation and analysis within the Taylor framework.
Furthermore, the additional experimental results on non-differentiable models (and the details of the models) are shown in Appendix C.
For the classification tasks in this study, we explain the model predictions based on the predicted probability of the positive class (label=1) rather than the final classification result (e.g., top-1 label). That is, the analysis is conducted with respect to the model’s continuous output — the estimated probability — rather than its discrete decision outcome.
Setup for fair comparison of TaylorPODA and the other baseline LA methods:
The proposed TaylorPODA, together with the existing OCC-1, SHAP, and WeightedSHAP, can be generalized as LA methods, thereby sharing similar calculation process. To establish fair comparison of TaylorPODA and the other LA baselines in the experiments, we further set up the corresponding parameters and designs by making full use of the similarity.
  • Since these LA methods rely on the masked outputs of the task models, they are configured to utilize a shared masked output calculator by incorporating one of the sub-functions weightedSHAP.generate_coalition_function provided within the WeightedSHAP package (https://github.com/ykwon0407/WeightedSHAP, used with the author’s permission).
  • According to (Kwon and Zou 2022) and the default settings of the WeightedSHAP package, WeightedSHAP generates solution candidates based on 16 distinct weight distributions. To ensure a relatively fair comparison, we similarly configure the search process of this version of TaylorPODA to include 16 distinct solution candidates. These candidates are generated from the Dirichlet distribution ($\alpha_{i}=1$ for all $i\in G$).
  • For the quantitative experimental results presented in Table 3, 5, and 6, we reimplemented a full version of SHAP without the subset sampling approximation, in accordance with (7), to avoid certain automatic approximation heuristics in the standard SHAP implementation. Similarly, we reimplemented a complete version of OCC-1 based on (8). This was done to faithfully follow the theoretical formulations and ensure consistency with the sharing use of the same masked outputs.
  • For the qualitative experimental results presented in Figure 4, we adopted SHAP (version 0.44.0, MIT license)’s PermutationExplainer to accommodate the use of approximation techniques. For LIME (version 0.2.0.1, BSD-2-Clause license), we utilized the LimeImageExplainer tailored for image data.
Appendix C C. Additional Experimental Results on Non-Differentiable Models
MLP with ReLU:
Table 5: Importance ordering performance on 100 hold-out test samples with MLP models and ReLU activation.
Classification Regression
Method
Data AUP Inclusion AUC Data AUP Inclusion MSE ($\times 10^{-2}$)
OCC-1
Cancer
1.042 (0.818, 1.267) 0.893 (0.855, 0.931)
Abalone
0.142 (0.124, 0.161) 0.067 (0.047, 0.088)
LIME 0.288 (0.199, 0.377) 0.989 (0.974, 1.000) 0.145 (0.128, 0.163) 0.057 (0.042, 0.072)
SHAP 0.241 (0.139, 0.343) 0.983 (0.967, 1.000) 0.171 (0.153, 0.189) 0.068 (0.053, 0.084)
WeightedSHAP 0.161 (0.110, 0.212) 0.991 (0.981, 1.000) 0.103 (0.091, 0.115) 0.034 (0.027, 0.042)
TaylorPODA 0.170 (0.106, 0.234) 0.991 (0.977, 1.000) 0.098 (0.085, 0.111) 0.030 (0.022, 0.038)
OCC-1
Rice
0.457 (0.359, 0.555) 0.951 (0.926, 0.977)
California
0.177 (0.152, 0.202) 0.137 (0.101, 0.173)
LIME 0.111 (0.076, 0.146) 0.997 (0.993, 1.000) 0.269 (0.238, 0.301) 0.304 (0.230, 0.377)
SHAP 0.118 (0.080, 0.156) 0.986 (0.973, 0.998) 0.190 (0.165, 0.215) 0.142 (0.105, 0.179)
WeightedSHAP 0.093 (0.062, 0.124) 0.991 (0.981, 1.000) 0.141 (0.121, 0.161) 0.097 (0.069, 0.125)
TaylorPODA 0.068 (0.048, 0.088) 0.997 (0.993, 1.000) 0.158 (0.135, 0.182) 0.111 (0.078, 0.143)
OCC-1
Titanic
0.330 (0.293, 0.367) 0.996 (0.990, 1.000)
Concrete
0.362 (0.312, 0.412) 0.551 (0.356, 0.747)
LIME 0.630 (0.535, 0.724) 0.944 (0.912, 0.977) 0.356 (0.323, 0.378) 0.388 (0.320, 0.456)
SHAP 0.483 (0.436, 0.530) 0.977 (0.954, 1.000) 0.273 (0.247, 0.299) 0.260 (0.215, 0.305)
WeightedSHAP 0.322 (0.285, 0.359) 0.997 (0.992, 1.000) 0.223 (0.201, 0.244) 0.199 (0.159, 0.238)
TaylorPODA 0.404 (0.367, 0.442) 0.989 (0.974, 1.000) 0.226 (0.203, 0.249) 0.204 (0.164, 0.244)
Figure 5: Discrepancy performance on 100 hold-out test samples with MLP models and ReLU activation.
XGBoost:
Table 6: Importance ordering performance on XGBoost (Chen and Guestrin 2016) models.
Classification Regression
Method
Data AUP Inclusion AUC Data AUP Inclusion MSE ($\times 10^{-2}$)
OCC-1
Cancer
0.600 (0.438, 0.762) 0.948 (0.922, 0.973)
Abalone
0.209 (0.182, 0.236) 0.117 (0.088, 0.147)
LIME 0.382 (0.301, 0.464) 0.992 (0.984, 1.000) 0.235 (0.207, 0.263) 0.131 (0.097, 0.165)
SHAP 0.280 (0.219, 0.340) 0.998 (0.995, 1.000) 0.226 (0.203, 0.248) 0.107 (0.084, 0.130)
WeightedSHAP 0.220 (0.189, 0.251) 1.000 (1.000, 1.000) 0.154 (0.131, 0.176) 0.066 (0.045, 0.088)
TaylorPODA 0.252 (0.209, 0.296) 1.000 (1.000, 1.000) 0.154 (0.131, 0.176) 0.065 (0.042, 0.087)
OCC-1
Rice
0.528 (0.425, 0.632) 0.951 (0.924, 0.979)
California
0.280 (0.234, 0.326) 0.289 (0.198, 0.381)
LIME 0.389 (0.265, 0.514) 0.971 (0.944, 0.999) 0.412 (0.353, 0.470) 0.547 (0.394, 0.699)
SHAP 0.325 (0.212, 0.438) 0.973 (0.946, 1.000) 0.362 (0.309, 0.415) 0.421 (0.293, 0.548)
WeightedSHAP 0.219 (0.156, 0.282) 0.976 (0.952, 0.999) 0.246 (0.202, 0.289) 0.240 (0.154, 0.326)
TaylorPODA 0.249 (0.169, 0.329) 0.983 (0.963, 1.000) 0.310 (0.257, 0.362) 0.347 (0.230, 0.464)
OCC-1
Titanic
0.584 (0.496, 0.671) 0.953 (0.917, 0.989)
Concrete
0.547 (0.475, 0.619) 0.962 (0.719, 1.205)
LIME 1.132 (0.952, 1.000) 0.877 (0.828, 0.926) 0.696 (0.617, 0.775) 1.328 (1.072, 1.584)
SHAP 0.916 (0.783, 1.000) 0.933 (0.896, 0.969) 0.574 (0.509, 0.639) 0.904 (0.724, 1.083)
WeightedSHAP 0.541 (0.459, 0.623) 0.957 (0.923, 0.991) 0.422 (0.361, 0.483) 0.602 (0.448, 0.757)
TaylorPODA 0.755 (0.654, 0.857) 0.950 (0.916, 0.984) 0.477 (0.411, 0.543) 0.732 (0.559, 0.905)
Figure 6: Discrepancy performance on 100 hold-out test samples with XGBoost models.
Appendix D D. Heuristic Approximation of TaylorPODA and its Error Bound
Although TaylorPODA offers a theoretically principled attribution method based on the full Taylor expansion, its direct application to high-dimensional data becomes infeasible due to the combinatorial explosion of high-cardinality interaction terms. As an initial step toward improving scalability and showcasing the potential of TaylorPODA on high-dimensional datasets, we propose a heuristic approximation (TaylorPODA-$c$) that preserves only low-cardinality terms over a restricted subset of input features:
[…]
where $c\in G$ with $c>1$ denotes a cap on the cardinality of $S$. Accordingly, as $c=2,\dots,|G|,\quad a_{i}^{(c)}(\mathbf{x},f(\mathbf{x}))\to a_{i}(\mathbf{x},f(\mathbf{x}))$. With the determined ceiling limit $c$, the approximation error between the attribution results produced by TaylorPODA-c and the full TaylorPODA is:
$\Delta_{i}(c;\mathbf{x},f)$ $=\Bigg{|}\underset{(\text{TaylorPODA-c})}{a_{i}^{(c)}(\mathbf{x},f(\mathbf{x}))}-\underset{(\text{TaylorPODA})}{a_{i}(\mathbf{x},f(\mathbf{x}))}\Bigg{|}$ (32)
$=\Bigg{|}\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>1\\ i\in S\end{subarray}}(1-\xi_{i,S})H(S)-\sum_{\begin{subarray}{c}S\subseteq G\\ 1<|S|\leq c\\ i\in S\end{subarray}}(1-\xi_{i,S})H(S)\Bigg{|}$
$=\Bigg{|}\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>c\\ i\in S\end{subarray}}(1-\xi_{i,S})H(S)\Bigg{|}.$
By capping the cardinality of the features, this approximation approach aligns with the insights provided by existing studies (Li and Zhang 2023; Ren et al. 2024) on the symbolic representations encoded by trained DNNs. Specifically, the Harsanyi dividend $H(S)$, referred to as “interaction” in (Li and Zhang 2023; Ren et al. 2024), has been shown to exhibit sparse value patterns (under some conditions that regulate a “well-trained” DNN)—“the DNN will only encode a relatively small number of sparse interactions between input variables”—which means that most $H(S)$ values are nearly zero. In particular, Li and Zhang (2023); Ren et al. (2024) have consistently shown that $|H(S)|$ remains below an empirical threshold of $0.05\cdot\max_{S^{\prime}}|H(S^{\prime})|$ (written as $0.05\bar{h}$) for most $S\subseteq G$ with $|S|>1$, especially when $|S|\gg 1$. That is, even relatively low-order interactions suffice to capture the majority of attribution mass, and higher-order interactions contribute only marginally.
Figure 7: Illustrative example of value distribution of $|H(S)|$, using the California dataset and the MLP model from previous experiments in this paper. The example shows consistent patterns with the results in (Ren et al. 2024)–most $|H(S)|$ remain below $0.05\bar{h}$. They indicate the rationality of truncating the higher-order interactions to approximate full TaylorPODA results.
This is further validated and exemplified in Figure 7, using the California dataset and the MLP model from previous experiments in this paper, where most interactions are non-salient. They exhibit almost zero contribution, with $|H(S)|\leq 0.05\bar{h}$ in the majority of cases.
Quantitatively, assuming that $f(\mathbf{x})$ is $(K+1)$ times differentiable and thus Taylor-expandable to order $(K+1)$, there exists some $\boldsymbol{\varrho}\in\mathbb{R}^{d}$ lying on the line segment between $\mathbf{x}$ and $\boldsymbol{\beta}$ such that (1) can be rewritten as
$f(\mathbf{x})$ $=f(\boldsymbol{\beta})+\sum_{i\in G}\frac{1}{1!}\cdot\frac{\partial f(\boldsymbol{\beta})}{\partial x_{i}}\cdot(x_{i}-\beta_{i})+\sum_{i\in G}\sum_{j\in G}\frac{1}{2!}\cdot\frac{\partial^{2}f(\boldsymbol{\beta})}{\partial x_{i}\partial x_{j}}\cdot(x_{i}-\beta_{i})(x_{j}-\beta_{j})$ (33)
$+\dots+\sum_{i\in G}\dots\sum_{j\in G}\frac{1}{(K+1))!}\cdot\frac{\partial^{K+1}f(\boldsymbol{\varrho})}{\partial x_{i}\dots\partial x_{j}}\cdot(x_{i}-\beta_{i})\cdots(x_{j}-\beta_{j}).$
[…]
$R_{f_{G\backslash\{i\}},\mathbf{x};\boldsymbol{\beta}}^{(K)}=\sum_{p\in G\backslash\{i\}}\dots\sum_{q\in G\backslash\{i\}}\frac{1}{(K+1))!}\cdot\frac{\partial^{K+1}f(\boldsymbol{\varrho})}{\partial x_{p}\dots\partial x_{q}}\cdot(x_{p}-\beta_{p})\cdots(x_{q}-\beta_{q}).$ (34)
Then we have
$R_{f,\mathbf{x};\boldsymbol{\beta}}^{(c)}-R_{f_{G\backslash\{i\}},\mathbf{x};\boldsymbol{\beta}}^{(c)}=\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>c\\ i\in S\end{subarray}}\sum_{\begin{subarray}{c}\psi\in\Pi_{S}\end{subarray}}\mu(\psi).$ (35)
Continuing from (32), we obtain an upper bound for $\Delta_{i}(c;\mathbf{x},f)$:
$\Delta_{i}(c;\mathbf{x},f)$ $=\Bigg{|}\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>c\\ i\in S\end{subarray}}(1-\xi_{i,S})H(S)\Bigg{|}\leq\Bigg{|}\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>c\\ i\in S\end{subarray}}H(S)\Bigg{|}=\Bigg{|}\sum_{\begin{subarray}{c}S\subseteq G\\ |S|>c\\ i\in S\end{subarray}}\sum_{\begin{subarray}{c}\psi\in\Pi_{S}\end{subarray}}\mu(\psi)\Bigg{|}$ (36)
[…]
where $M$ denotes an upper bound on the absolute value of all $(c+1)$-th order partial derivatives of $f$ within the region defining all the possible $\boldsymbol{\varrho}$’s, i.e.,
$M=\max_{\boldsymbol{\varrho}\in[\mathbf{x},\boldsymbol{\beta}]}\max_{|\alpha|=c+1}\left|\frac{\partial^{|\alpha|}f(\boldsymbol{\varrho})}{\partial\mathbf{x}^{\alpha}}\right|.$ (37)
In practice, the upper bound of $\Delta_{i}(c;\mathbf{x},f)$ given in (36) is often small in magnitude, especially when the evaluation point $\mathbf{x}$ lies sufficiently close to the baseline $\boldsymbol{\beta}$, or when the function $f$ exhibits limited higher-order variability. This is commonly observed in the following scenarios:
  • When $f$ is a low-degree polynomial (e.g., quadratic or cubic), the remainder term $R_{f,\mathbf{x};\boldsymbol{\beta}}^{(c)}$ vanishes for all $c$ greater than the polynomial degree. For example, if $f(\mathbf{x})$ is quadratic, then $\Delta_{i}(c;\mathbf{x},f)=0$ for all $c\geq 2$.
  • When $f$ is a smooth neural network with activation functions like tanh or sigmoid, the higher-order partial derivatives tend to decay rapidly, especially when $\|\mathbf{x}-\boldsymbol{\beta}\|$ is small due to normalization or local linearity.
  • In kernel methods, such as those using RBF kernels, the function $f(\mathbf{x})$ is often extremely smooth (infinitely differentiable), and higher-order derivatives are exponentially suppressed with respect to the distance between $\mathbf{x}$ and $\boldsymbol{\beta}$.
Therefore, in many practical settings, especially in locally smooth regions or when $c$ is moderately large, the quantity $\Delta_{i}(c;\mathbf{x},f)$ remains negligible and contributes little to the overall approximation error. While more sophisticated and broadly applicable approximation methods remain an open direction for future work, we believe this heuristic provides a reasonable and principled starting point.